OpenStudy (superhelp101):

1. A discriminant of 0 indicates what type of roots? Two Irrational Two Complex Two Rational One Rational Double 2. The quadratic x2 - 7x + 6 has what type of roots? Two Irrational Two Complex Two Rational One Rational Double 3. The quadratic x2 + 5x - 14 has what type of roots? Two Irrational Two Complex Two Rational One Rational Double 4. What is the possible discriminant of the graph? a parabola opening up with vertex near negative 1.2 comma negative 9.1. a y intercept of negative 6 -11 zero 25 73 5. What is the possible discriminant of the graph? a parabola opening up x intercepts near negative 7 and negative 3 -13 zero 15 16

OpenStudy (superhelp101):

I got D, C, C, D. Just make sure those are right. And don't get the last one too. @ganeshie8

OpenStudy (superhelp101):

just making sure*

ganeshie8 (ganeshie8):

1, 2, 3 questions are correct im not sure about 4 and 5

ganeshie8 (ganeshie8):

I'll need to check the graph to confirm questions #4 and #5

OpenStudy (superhelp101):

oh okay. I will attach them, just a sec...

OpenStudy (superhelp101):

@ganeshie8

ganeshie8 (ganeshie8):

@Kainui

OpenStudy (anonymous):

1)D one rational double because when the discriminant is zero you get one answer 2) C 2 rational roots because b^2 - 4ac a = 1, b = - 7, c = 6 and 49-4(1)(6)=25 since 25 is greater than zero it has 2 rational roots

OpenStudy (anonymous):

3)C 2 rational roots. x=2,-7 because of (x-2)(x+7)

OpenStudy (superhelp101):

Awww thanks! do you understand #4 and #5

OpenStudy (anonymous):

4) no not really 5) no not really sorry:(

OpenStudy (superhelp101):

awww it's ok